OpenStudy (anonymous):

Solve for x/8 = 3 3/4 . A. 24 3/4 B. 15/32 C. 120 D. 30

OpenStudy (sleepyhead314):

is that a \(3 \frac{ 3 }{ 4 }\) ?

OpenStudy (anonymous):

yes..............how did you do that?

OpenStudy (sleepyhead314):

the blue "equation" button below :)

OpenStudy (sleepyhead314):

now, first let's convert that into an improper fraction can you do that? :)

OpenStudy (anonymous):

x 3 with 4 and add 3

OpenStudy (anonymous):

15/ 4

OpenStudy (sleepyhead314):

good :) so now we have\[\frac{ x }{ 8 }= \frac{ 15 }{ 4 }\]multiply both sides by 8, what would you get?

OpenStudy (anonymous):

So 8 times 8?

OpenStudy (sleepyhead314):

oh (1/8) times 8 = 1

OpenStudy (anonymous):

what?

OpenStudy (sleepyhead314):

\[\frac{ x }{ 8 }*\frac{ 8 }{ 1 }=x\]

OpenStudy (sleepyhead314):

it's just something that is true and if we multiply one side by 8 then we have to multiply the other side too \[x = \frac{ 15 }{ 4 }*\frac{ 8 }{ 1 }=~?\]do you know how to multiply fractions? :)

OpenStudy (anonymous):

15 x 8 = 120 4 x 1 = 4

OpenStudy (anonymous):

Oh so ur saying that 15 is the answer?

OpenStudy (sleepyhead314):

ehhh 120 divided by 4 = ?

OpenStudy (anonymous):

120 divided by 4 = 30

OpenStudy (sleepyhead314):

yes :)

OpenStudy (anonymous):

wait so 30 is the answer not 15.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

30 is the answer.

OpenStudy (anonymous):

next question.